# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v3

@@ 4.3. Outlier Value Encoding through N:M Structured Pruning
To formalize step 2, let us assume $n$ outliers are present in a $\mu$B quantized to MX-FP-44448,8. After sharing 1-bit $\mu X$ across the $n$ outliers, each outlier has 1 sign ($s$) and 2 mantissa bits ($m_{1}m_{0}$). The inliers are 2-bit 2’s complement MX-INT and has 1-bit sign and 1-bit magnitude. Since the outlier bits have two mantissa bits with one sign, to ensure symmetric outlier distribution, we duplicate their sign bit and assign each sign to a mantissa creating and partitioning into two halves Upper, Lower of size 2-bit each mimicking inlier MX-INT structure i.e., $\{sm_{1},sm_{0}\}$ (see Figure 3(a), step 2). While we demonstrate this for a specific example, the process can be generalized to other inlier and outlier formats bit-widths, wherein the remaining outlier bits need to be reorganized into Upper and Lower distributable halves each of size $bb$ (per element bit-budget), with $bb$ representing the operational bit-width of each PE.
In step 3, to ensure a fixed $bb$ and data-type of a layer for aligned memory access and simple PE design, we distribute the outlier LSBs (Lower half) to the least important inlier locations that are pruned within a $\mu$B. For $n$ outliers in a $\mu$B we identify $n$ least important inlier locations to prune via the Hessian information (see Algorithm 1), forming a ($B_{\mu}$-$n$):$B_{\mu}$ structured pruning pattern (Jeong et al., 2023; Sun et al., 2021) i.e., ($B_{\mu}$-$n$) non-zero values exist per $\mu$B after pruning. To keep track of the corresponding halves of each of the outliers in a $\mu$B, we maintain a per-$\mu$B permutation list (Note: if there are no outliers in a $\mu$B, a permutation list is not stored). The permutation list for each $\mu$B that has outliers is made up of $\frac{B_{\mu}}{2}$ elements (maximum number of outliers supported per $\mu B$) with each element storing the locations of the Upper and Lower halves of outliers in a 6-bit format $\{\texttt{Upper}_{loc},\texttt{Lower}_{loc}\}$ (see Figure 3(a) step 3). Note that, in case of outlier count $>\frac{B_{\mu}}{2}$, the mB size should be chosen to be higher to prevent pruning of outliers and causing higher accuracy degradation. While such a situation does not arise in any of the models that we evaluate with $B_{\mu}=8$, this feature only serves as a demonstration of the flexibility of MicroScopiQ for future models.
@@ 5.1. Architecture Overview
In Figure 4, we depict the MicroScopiQ accelerator architecture, integrated into a standard weight stationary systolic array. The accelerator is optimized for throughput and all compute units (PE array, ReCoN etc.) are internally pipelined with interleaved pipeline stages. Through the numbered steps in Figure 4, we explain the computational flow of the MicroScopiQ accelerator. And in Figure 5(a-c), we demonstrate the mapping and schedule for a sample $2\times 2$ GEMM operation with inlier and outlier weights, following the same computation flow.
@@
PE row 1 receives partial sums from the top and similarly performs INT multiplication between weights and iAct. As shown in the scheduling figure, in parallel, the output scale factors can be calculated (section 5.5). 4 The mapped $\mu$Bs to PE row 1 have outliers, the controller directs all PE outputs to ReCoN. PE row 1 offloads the accumulation to ReCoN (shown by dotted lines in Figure 5(c)). This is needed because the outlier will have its two halves distributed in different locations and mapped to different PEs. While the PEs in row 0 with inlier weights can accumulate the partial sum output, the PEs in row 1 with the Upper and Lower outlier halves cannot compute the outlier partial sum output. After receiving outputs from row 1, ReCoN reorders and calculates the FP-outlier partial sum.
@@ 5.2. On-chip Storage and Control
Weight/Activation Organization. The iActs are stored in the iAct buffer as 8-bit INTs. Lower-precision iActs ($<$8-bit) are supported through sign-extension. The same process is followed for oActs. The weight buffer stores weights at 4-bit granularity. At lower precision (2-bits), each buffer location simultaneously stores two 2-bit weights. The MODE signal from the controller identifies the interpretation of bit format of weights in the weight buffer i.e, one 4-bit weight or two 2-bit weights.
@@
MicroScopiQ Controller. The controller generates appropriate control and signals to all units. It is functionally very similar to standard systolic array controllers (Sharma et al., 2018), however, with added functionality to support specific features of ReCoN (arbitration of simultaneous access by multiple rows), multi-precision PEs, post-processing unit. Furthermore, unlike traditional controllers, the MicroScopiQ controller exerts fine-grained control using handshaking signals on the streaming of iActs, iAccs into the PE array to account for the increased pipeline depth through ReCoN for rows with outliers.
@@ 5.4. Redistribution and Coordination NoC
ReCoN Input Interface. ReCoN is shared and time-multiplexed across all PE rows, and at a particular instant many rows might require access to it. At the inputs, the column-wise arbiters guided by control signals from the controller, coordinate and resolve contention between different PE rows and ensure a fair scheduling (Note for simplicity in Figure 4, we conceptually depict this operation as a N-input multiplexer). Due to the skewed data flow of systolic array computations (Jouppi et al., 2017; Ramachandran et al., 2024b), partial sums from different columns of the same row do not arrive at the ReCoN input at the same cycle. A synchronization buffer (Figure 6(c)) is positioned between ReCoN and column-wise multiplexers which propagates the partial sums through sets of buffers of different lengths (column producing the partial sum fastest has the most buffers), synchronizing the arrival of inputs from different columns. The synchronization buffer sends an ACK signal to the PE row whose values have been accepted so that it can progress in its computation and the other simultaneously contesting rows will be on hold and acknowledged in the next cycle. When there are N simultaneously contesting rows, the row that is acknowledged the last adds an N-1 cycle latency to its processing. As we show in Figure 10(b), Figure 13(a) the num. of conflicts in accessing ReCoN are minimal and if latency is a critical factor multiple ReCoN units can be used to minimize conflicts.
@@
5.5. Post Processing and Output Scale Factor Computation
@@
Table 2. Quantization Results for LLMs. We report WikiText2 (Merity et al., 2016) perplexity numbers (lower the better).
@@
Table 3. LLM quantization performance of MicroScopiQ augmented with OmniQuant (Omni-MicroScopiQ).
  Method W/A Llama-2 13B Llama-3 70B Phi-3 3.8B
Baseline 16/16 4.83 2.85 6.33
  OmniQuant (Shao et al., 2023) 4/16 5.02 3.46 6.67
Omni-MicroScopiQ 4/16 4.87 2.97 6.52
  OmniQuant (Shao et al., 2023) 2/16 7.56 6.17 7.09
Omni-MicroScopiQ 2/16 6.58 5.09 6.89
  OmniQuant (Shao et al., 2023) 2/8 8.92 6.83 7.95
Omni-MicroScopiQ 2/8 7.12 5.74 7.21
  Method W/A ARC-c (Clark et al., 2018) HellaS. (Zellers et al., 2019) MMLU (Hendrycks et al., 2020) WinoG. (Sakaguchi et al., 2021)
Baseline 16/16 60.50 84.30 68.90 80.60
  OliVe (Guo et al., 2023) 2/16 38.60 55.30 39.80 60.70
OmniQuant (Shao et al., 2023) 2/16 49.70 77.80 58.20 74.20
MicroScopiQ 2/16 53.30 81.60 63.70 77.80
@@ 5.6. MicroScopiQ Walkthrough Example
The $\mu$B after distributing the outlier is mapped to PE row 0. 3 PE row 0 computes partial sum outputs for inliers in columns 1 and 2, while it outputs $O_{\texttt{Upper},\texttt{Res}}$ and $O_{\texttt{Lower},\texttt{Res}}$ for the other two columns that have outliers. The output of PE row 0 is routed through ReCoN. In ReCoN, the inliers are passed down at all stages, since they are already present in their respective columns.
@@
The stage 1 switch at column 3, executes a swap operation and redirects the lower half outlier output towards the upper half while simultaneously sending iAcc through the other port. This iAcc corresponds to the iAcc input to column 3 PE in row 0. This is functionally correct because, the lower half of outlier is distributed in this column by zeroing out/pruning the inlier at that location, hence the partial sum output of that PE to the corresponding PE in the next row should only be iAcc since the original weight at that location is zero. Similarly in stage 2 of ReCoN, the switch that receives the $O_{\texttt{Lower},\texttt{Res}}$ from stage 1 again executes a swap operation. Finally in stage 3 the outlier output halves are merged together to realize $O^{0}_{Res}$, while all other switches execute pass. This merge operation results in a $O^{0}_{Res}$ that matches the expected output.
@@
The partial sums from ReCoN are then sent to PE row 1 which then perform MAC operations and sends it to the subsequent row since there are no outlier weights. Similarly all rows execute operations until the final partial sum is calculated.
@@ 6. Experimental Evaluations
Models and Datasets. We evaluate on OPT (Zhang et al., 2022), LLaMA2 (Touvron et al., 2023) LLaMA3 (Meta, 2024), Phi-3 (SLM) (Abdin et al., 2024) and Mixtral (MoE) (Jiang et al., 2024) LLM model families. The VLMs evaluated are OpenFlamingo (Awadalla et al., 2023), VILA (Lin et al., 2024b) and LlaVa (Liu et al., 2024a). We use a calibration dataset consisting of 256 random samples from the PILE dataset (Gao et al., 2020), so as to not overfit to any particular downstream domain (Lin et al., 2024a). We compare different LLMs based on perplexity (PPL $\downarrow$) on the WikiText2 (Merity et al., 2016) dataset and benchmark accuracy on 6 different tasks BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-c (Clark et al., 2018), MMLU (Hendrycks et al., 2020) and Winogrande (Sakaguchi et al., 2021). Similarly, we evaluate VLMs on 5 vision-language benchmarks: COCO captioning (Chen et al., 2015), VQAv2 (Goyal et al., 2017), VizWiz (Gurari et al., 2018), TextVQA (Singh et al., 2019), and GQA (Hudson and Manning, 2019).
@@
Accelerator Implementation. The MicroScopiQ accelerator is implemented in Verilog RTL. We perform synthesis and PnR using Synopsys Design Compiler and Innovus with a TSMC 7nm technology library (See floorplan in Figure 11(b)) for area, power and latency calculations. All MicroscopiQ design variants attains a peak clock frequency of 1 GHz. We use CACTI (Muralimanohar et al., 2009) to estimate the area and power of on-chip memories. For end-to-end performance evaluation, we design a cycle accurate simulator based on DnnWeaver (Sharma et al., 2016) and BitFusion (Sharma et al., 2018), following previous works (Guo et al., 2023; Ramachandran et al., 2024b).
Accelerator Baselines. We compare the MicroScopiQ accelerator against OliVe (Guo et al., 2023), ANT (Guo et al., 2022), GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) and AdaptivFloat (Tambe et al., 2020) across area, power and performance metrics. To demonstrate the minimal overhead of ReCoN on NoC-based real-world industrial and academic accelerators, we also implement and compare against baseline MTIA (Firoozshahian et al., 2023) and Eyeriss-v2 (Chen et al., 2019). For a fair comparison, we ensure all designs attain a clock frequency of 1GHz. DeepScale (Sarangi and Baas, 2021) is employed to scale all designs to the 7nm process.

### v4

@@ 4.3. Outlier Value Encoding through N:M Structured Pruning
To formalize Step 2, let us assume $n$ outliers are present in a $\mu$B quantized to MX-FP-44448,8. After sharing 1-bit $\mu X$ across the $n$ outliers, each outlier has 1 sign ($s$) and 2 mantissa bits ($m_{1}m_{0}$). The inliers are 2-bit 2’s complement MX-INT and has 1-bit sign and 1-bit magnitude. Since the outlier bits have two mantissa bits with one sign, to ensure symmetric outlier distribution, we duplicate their sign bit and assign each sign to a mantissa creating and partitioning into two halves Upper,Lower of size 2-bit each mimicking inlier MX-INT structure i.e., $\{sm_{1},sm_{0}\}$ (see Figure 3(a), Step 2). While we demonstrate this for a specific example, the process can be generalized to other inlier and outlier formats bit-widths wherein, the remaining outlier bits need to be reorganized into Upper and Lower distributable halves each of size $bb$ (per element bit-budget), with $bb$ representing the operational bit-width of each PE.
In Step 3, to ensure a fixed $bb$ and data-type of a layer for aligned memory access and simple PE design, we distribute the outlier LSBs (Lower half) to the least important inlier locations that are pruned within a $\mu$B. For $n$ outliers in a $\mu$B we identify $n$ least important inlier locations to prune via the Hessian information (see Algorithm 1), forming a ($B_{\mu}$-$n$):$B_{\mu}$ structured pruning pattern  [^3] Following the structure of a standard N:M pattern (Jeong et al., 2023) where N = ($B_{\mu}$-$n$) and M = $B_{\mu}$ (Jeong et al., 2023; Sun et al., 2021) i.e., ($B_{\mu}$-$n$) non-zero values exist for every $B_{\mu}$ after pruning. To keep track of the corresponding halves of each of the outliers in a $\mu$B, we maintain a per-$\mu$B permutation list (Note: if there are no outliers in a $\mu$B, a permutation list is not stored). The permutation list for each $\mu$B that has outliers is made up of $\frac{B_{\mu}}{2}$ elements (maximum number of outliers supported per $\mu B$) with each element storing the locations of the Upper and Lower halves of outliers in a 6-bit format $\{\texttt{Upper}_{loc},\texttt{Lower}_{loc}\}$ (see Figure 3(a) Step 3). Note that, in case of outlier count $>\frac{B_{\mu}}{2}$, the mB size should be chosen to be higher to prevent pruning of outliers and causing higher accuracy degradation. While such a situation does not arise in any of the models that we evaluate with $B_{\mu}=8$, this feature only serves as a demonstration of the flexibility of MicroScopiQ for future models.
@@ 5.1. Architecture Overview
In Figure 4, we depict the MicroScopiQ accelerator architecture, integrated into a standard weight stationary systolic array. The accelerator is optimized for throughput and all compute units (PE array, ReCoN etc.) are internally pipelined with interleaved pipeline stages. Figure 5 illustrates MicroScopiQ’s interaction with on-chip and off-chip memories. Following prior work (Keller et al., 2023; Fang et al., 2024; Dai et al., 2021), off-chip memory is modeled as HBM2 with a 256 GB/s bandwidth. The off-chip memory layout (Guo et al., 2023; Zadeh et al., 2020) of a MicroScopiQ-quantized tensor, shown in Figure 5, consists of two sections: quantized weights and metadata. Following prior work (Guo et al., 2023, 2022; Ramachandran et al., 2024b) a two-level on-chip memory hierarchy is employed: 2MB L2 global SRAM loads data from DRAM and transfers it to MicroScopiQ buffers (section 5.2) via an Open Compute Protocol OCP-SRAM interface (Sumbul et al., 2022) with a bandwidth of 64 GB/s.
Through the numbered steps in Figure 4, we explain the computational flow of the MicroScopiQ accelerator. And in Figure 6(a-c), we demonstrate the mapping and schedule for a sample $2\times 2$ GEMM operation with inlier and outlier weights, following the same computation flow.
@@
PE row 1 receives partial sums from the top and similarly performs INT multiplication between weights and iAct. As shown in the scheduling figure, in parallel, the output scale factors can be calculated (section 5.5).
4
The mapped $\mu$Bs to PE row 1 have outliers, the controller directs all PE outputs to ReCoN. PE row 1 offloads the accumulation to ReCoN (shown by dotted lines in Figure 6(c)). This is needed because the outlier will have its two halves distributed in different locations and mapped to different PEs. While the PEs in row 0 with inlier weights can accumulate the partial sum output, the PEs in row 1 with the Upper and Lower outlier halves cannot compute the outlier partial sum output. After receiving outputs from row 1, ReCoN reorders and calculates FP-outlier partial sum.
@@ 5.2. On-chip Storage and Control
Weight/Activation Organization. The input activations (iActs) are stored in the iAct buffer as 8-bit INTs. Lower-precision iActs ($<$8-bit) are supported through sign-extension. The same process is followed for oActs (output activations). The weight buffer stores weights at 4-bit granularity. At lower precision (2-bits), each buffer location simultaneously stores two 2-bit weights. The MODE signal from the controller delineates the bit format of weights in the weight buffer i.e, one 4-bit weight or two 2-bit weights.
@@
MicroScopiQ Controller. The controller generates appropriate control and signals to all units. It is functionally very similar to standard systolic array controllers (Sharma et al., 2018), however, with added functionality to support specific features of ReCoN (arbitration of simultaneous access by multiple rows), multi-precision PEs, post-processing unit. Furthermore, unlike traditional controllers, the MicroScopiQ controller exerts fine-grained control using handshaking signals on the streaming of iActs, iAccs into the PE array to account for the increased pipeline depth through ReCoN for rows with outliers. The controller also generates OCP-SRAM interface control signals to manage data flow between the buffers and the L2 global SRAM.
@@ 5.4. Redistribution and Coordination NoC
ReCoN Input Interface. ReCoN is shared and time-multiplexed across all PE rows, and at a particular instant many rows might require access to it. At the inputs, the column-wise arbiters guided by control signals from the controller, coordinate and resolve contention between different PE rows and ensure a fair scheduling (for simplicity in Figure 4, we conceptually depict this operation as an N-input multiplexer). Due to the skewed data flow of systolic array computations (Jouppi et al., 2017; Ramachandran et al., 2024b), partial sums from different columns of the same row do not arrive at the ReCoN input at the same cycle. A synchronization buffer (Figure 7(c)) is positioned between ReCoN and column-wise multiplexers which propagates the partial sums through sets of buffers of different lengths (column producing the partial sum fastest has the most buffers), synchronizing the arrival of inputs from different columns. The synchronization buffer sends an ACK signal to the PE row whose values have been accepted so that it can progress in its computation and the other simultaneously contesting rows will be on hold and acknowledged in the next cycle. When there are N simultaneously contesting rows, the row that is acknowledged the last adds an N-1 cycle latency to its processing. As we show in Figure 14(b), Figure 18(a) the num. of conflicts in accessing ReCoN are minimal and if latency is a critical factor multiple ReCoN units can be used to minimize conflicts.
@@
5.5. Post Processing and Scale Compute Unit
@@
Figure 9. Illustration of W4A4 MicroScopiQ GEMM on a GPU.
Table 2. Quantization Results for LLMs. We report WikiText2 perplexity numbers (lower the better).
@@ 5.6. MicroScopiQ Walkthrough Example
The $\mu$B after distributing the outlier is mapped to PE row 0.
3
PE row 0 computes partial sum outputs for inliers in columns 1 and 2, while it outputs $O_{\texttt{Upper},\texttt{Res}}$ and $O_{\texttt{Lower},\texttt{Res}}$ for the other two columns that have outliers. The output of PE row 0 is routed through ReCoN. In ReCoN, the inliers are passed down at all stages, since they are already present in their respective columns.
@@
The level 1 switch at column 3, executes a swap operation and redirects the lower half outlier output towards the upper half while simultaneously sending iAcc through the other port. This iAcc corresponds to the iAcc input to column 3 PE in row 0. This is functionally correct because, the lower half of outlier is distributed in this column by zeroing out/pruning the inlier at that location, hence the partial sum output of that PE to the corresponding PE in the next row should only be iAcc since the original weight at that location is zero. Similarly in level 2 of ReCoN, the switch that receives the $O_{\texttt{Lower},\texttt{Res}}$ from level 1 again executes a swap operation. Finally in level 3 the outlier output halves are merged together to realize $O^{0}_{Res}$, while all other switches execute pass. This merge operation results in a $O^{0}_{Res}$ that matches the expected output.
@@
, 6 The partial sums from ReCoN are then sent to PE row 1 which then perform MAC operations and sends it to the subsequent row since there are no outlier weights. Similarly all rows execute operations until the final partial sum is calculated.
6. MicroScopiQ Integration in GPUs
In this section, we demonstrate the integration of MicroScopiQ into GPUs via SW (section 6.1) and HW (section 6.2) modifications. For a matrix multiplication problem of size $M\times N\times K$, each GPU thread block (Lin et al., 2024b; Zhao et al., 2024) is responsible for computing a $T_{m}\times T_{n}$ output tile. This computation is performed iteratively along the $K$ dimension (Figure 9), which is referred to as the main loop. MicroScopiQ quantized models face acceleration challenges on GPUs due to: (a) high-cost pointer arithmetic from distributed outliers and (b) tensor cores’ inability to co-issue INT and FP operations (Corporation, 2017), preventing direct acceleration of tiles. Our SW, HW optimizations address these issues to maximize GPU performance.
6.1. SW: Kernel-Level Optimizations
Register Caching. We implement a virtual caching layer (Ben-Sasson et al., 2016) to efficiently distribute outliers within a warp. Instead of each thread block loading and merging outliers from shared memory, we use shfl_sync(m, r, t) (Ben-Sasson et al., 2016) for intra-warp communication. This primitive allows a thread to share its register r while reading the value from thread t within the same warp, using m as the thread selector mask. With $B_{\mu}=8$, each warp (32 threads) has 4 $\mu B$s, allowing efficient intra-warp outlier merging based on the permutation list.
GEMM on Tensor Cores. Due to MicroScopiQ’s MX-FP format for outliers and MX-INT for inliers, mixed tiles after outlier merging (iteration 0 in Figure 9) must first be dequantized to FP16 before GEMM execution. In contrast, inlier-only tiles (iteration 1) can leverage 4-bit INT tensor cores for efficient acceleration but still require INT32-to-FP16 dequantization for accumulation along the K-dimension (Figure 9) with other tiles. We implement this block-level dynamic decision to deliver maximum performance. However, the need for repeated dequantization and lack of multi-precision support is a significant bottleneck on GPUs.
6.2. HW: Potential Tensor Core Modification
Following prior modeling (Guo et al., 2023; Raihan et al., 2019), each tensor core (Corporation, 2017; Abdelkhalik et al., 2022) performs 16-bit FEDPs (four-element dot products). Therefore, each tensor core at 4-bits can conduct 16EDPs. Efficiently accelerating MicroScopiQ requires simultaneous INT+FP support within the tensor core. Following the functionality of ReCoN in section 5.4, each 16EDP operation requires a variable right shifter (Inliers: $>>0$, Outlier Upper Half: $>>1$, Outlier Lower Half: $>>2$) to account for the FP mantissa for outlier products. With typical GPU die sizes (RTX 2080 Ti: 754 $mm^{2}$), adding a shifter has negligible overhead ($\sim 0.1\%$).
@@ 7. Experimental Evaluations
Models and Datasets. We evaluate on OPT (Zhang et al., 2022), LLaMA2 (Touvron et al., 2023) LLaMA3 (Meta, 2024), Phi-3 (SLM) (Abdin et al., 2024) and Mixtral (MoE) (Jiang et al., 2024) LLM model families. The VLMs evaluated are OpenFlamingo (Awadalla et al., 2023), VILA (Lin et al., 2024c) and LlaVa (Liu et al., 2024a). We use a calibration dataset consisting of 256 random samples from the PILE dataset (Gao et al., 2020), so as to not overfit to any particular downstream domain (Lin et al., 2024a). We compare different LLMs based on perplexity (PPL $\downarrow$) on the WikiText2 (Merity et al., 2016) dataset and benchmark accuracy on 6 different tasks BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-c (Clark et al., 2018), MMLU (Hendrycks et al., 2020) and Winogrande (Sakaguchi et al., 2021).
Similarly, we evaluate VLMs on 5 vision-language benchmarks: COCO captioning (Chen et al., 2015), VQAv2 (Goyal et al., 2017), VizWiz (Gurari et al., 2018), TextVQA (Singh et al., 2019), and GQA (Hudson and Manning, 2019). Additionally, to demonstrate generalizability we also evaluate Convolutional Neural Network (CNN) baselines-ResNet50, VGG16 and State Space Model (SSM) baselines-VMamba (Liu et al., 2024b), Vim (Zhu et al., 2024). For these models we employ a calibration dataset of 64 samples from the Imagenet training set (Ramachandran et al., 2025b).
@@
Accelerator Implementation. The MicroScopiQ accelerator is implemented in Verilog RTL. We perform synthesis and PnR using Synopsys Design Compiler and Innovus with a TSMC 7nm technology library (See floorplan in Figure 16(a)) for area, power and latency calculations. All MicroscopiQ design variants attains a peak clock frequency of 1 GHz. We use CACTI (Muralimanohar et al., 2009) to estimate the area and power of on-chip memories. For end-to-end performance evaluation, we design a cycle accurate simulator based on DnnWeaver (Sharma et al., 2016) and BitFusion (Sharma et al., 2018), following previous works (Guo et al., 2023; Ramachandran et al., 2024b).
Accelerator Baselines. We compare the MicroScopiQ accelerator against OliVe (Guo et al., 2023), ANT (Guo et al., 2022), GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) and AdaptivFloat (Tambe et al., 2020) across area, power and performance metrics. To demonstrate the minimal overhead of ReCoN on NoC-based real-world industrial and academic accelerators, we also implement and compare against baseline MTIA (Firoozshahian et al., 2023) and Eyeriss-v2 (Chen et al., 2019). For a fair comparison, we ensure all designs attain a clock frequency of 1GHz and have same memory hierarchy and bandwidth. DeepScale (Sarangi and Baas, 2021) is employed to scale all designs to the 7nm process.
GPU Implementation. The optimized MicroScopiQ kernel is implemented in CUDA (Zhao et al., 2024) with a PyTorch frontend. We extend GPGPU-Sim (Bakhoda et al., 2009) and AccelSim (Khairy et al., 2018) for Ampere GPUs to integrate and evaluate the modified tensor core. Energy estimation is performed using AccelWattch (Khairy et al., 2020), GPUWattch (Leng et al., 2013). Real GPU inference throughput follows the setup in (Zhao et al., 2024), while simulated GPU results follow (Guo et al., 2023).
GPU Baselines. We compared system-level performance with FP16 TensorRT-LLM (NVIDIA, 2024) baseline and W4A4 Atom kernel (Zhao et al., 2024).
  Method W/A ARC-c HellaSwag MMLU WinoGrande
Baseline 16/16 60.50 84.30 68.90 80.60
  OliVe (Guo et al., 2023) 2/16 38.60 55.30 39.80 60.70
OmniQuant (Shao et al., 2023) 2/16 49.70 77.80 58.20 74.20
MicroScopiQ (Ours) 2/16 53.30 81.60 63.70 77.80